### v2

# EfficientNet with Hybrid Attention Mechanisms for Enhanced Breast Histopathology Classification: A Comprehensive Approach

Abstract
Breast cancer histopathology image classification is crucial for early cancer detection, offering the potential to reduce mortality rates through timely diagnosis. This paper introduces a novel approach integrating Hybrid EfficientNet models with advanced attention mechanisms, including Convolutional Block Attention Module (CBAM), Self-Attention, and Deformable Attention, to enhance feature extraction and focus on critical image regions. We evaluate the performance of our models across multiple magnification scales using publicly available histopathological datasets. Our method achieves significant improvements, with accuracy reaching 98.42% at 400X magnification, surpassing several state-of-the-art models, including VGG and ResNet architectures. The results are validated using metrics such as accuracy, F1-score, precision, and recall, demonstrating the clinical potential of our model in improving diagnostic accuracy. Furthermore, the proposed method shows increased computational efficiency, making it suitable for integration into real-time diagnostic workflows.
Index Terms:
Transfer Learning, Modality Specific, EfficientNet, CBAM, Self-Attention, Deformable Attention, Histopathology, Image Classification, Deep Learning, Breast Cancer.
I Introduction
Breast cancer remains one of the most prevalent and deadly forms of cancer, especially affecting women worldwide. According to global cancer statistics, breast cancer accounts for more than 2.3 million new cases each year, contributing to a significant number of cancer-related deaths [3]. Early detection plays a critical role in reducing the mortality rate associated with this disease. The standard diagnostic procedure for detecting breast cancer involves histopathological analysis, where biopsy tissue samples are examined under a microscope. This process is time-consuming, labor-intensive, and prone to subjective interpretation errors, even among trained pathologists [1].
Pathologists rely on morphological abnormalities in cell structures, particularly the nuclei, to distinguish between benign and malignant cells [2]. Benign tumors typically exhibit well-organized cellular structures and rarely invade surrounding tissues or metastasize to other parts of the body. In contrast, malignant tumors display uncontrolled growth, irregular nuclei, and the potential to spread, making them life-threatening. Early identification of these malignant traits is critical for improving patient prognosis and guiding treatment plans.
In recent years, computer-aided diagnosis (CAD) systems have been developed to assist pathologists by automating the analysis of histopathology images. Deep learning (DL) models, particularly Convolutional Neural Networks (CNNs), have shown great promise in this domain due to their ability to automatically learn complex features from medical images [7]. However, despite their success in natural image classification, traditional CNN architectures, such as VGGNet and ResNet, struggle to capture the intricate features present in histopathological images, primarily due to the variability in magnification scales and the complex structure of the tissue samples.
To address these challenges, researchers have begun exploring hybrid models that combine CNNs with attention mechanisms. These attention mechanisms, including the Convolutional Block Attention Module (CBAM), Self-Attention, and Deformable Attention, have been designed to enhance feature extraction by allowing the network to focus on the most relevant parts of the image. CBAM, for example, has shown considerable potential in medical image analysis by directing the network’s attention toward critical regions, leading to improved classification accuracy. Similarly, Self-Attention mechanisms, initially introduced in transformer models for Natural Language Processing (NLP), have been adapted to medical imaging tasks, where they capture long-range dependencies across the image. Deformable Attention further enhances this by dynamically adjusting the receptive field based on the complexity of the input features, making it particularly effective for identifying abnormalities in histopathology images.
In this study, we propose a series of Hybrid EfficientNet models integrated with advanced attention mechanisms to improve the performance of breast cancer histopathology image classification. EfficientNet, a highly scalable CNN architecture, has been recognized for its ability to achieve state-of-the-art performance with fewer parameters by optimizing the depth, width, and resolution of the model. By incorporating attention mechanisms such as CBAM, Self-Attention, and Deformable Attention into EfficientNet, we aim to enhance the network’s focus on key features, thereby improving the classification of benign and malignant breast cancer tissue samples.
The dataset used in this study is the BreakHis dataset, a publicly available dataset widely used for breast cancer histopathology image classification tasks . The dataset consists of 9,109 images at four different magnification levels (40X, 100X, 200X, and 400X), providing a comprehensive set of data to test the proposed models across varying scales of tissue detail. The performance of the proposed models is evaluated using key metrics such as accuracy, precision, recall, and F1-score.
The rest of the paper is structured as follows: In Section II, we review the existing literature on breast cancer histopathology image classification and attention mechanisms. Section III describes the methodology, including the dataset, preprocessing techniques, model architecture, and training procedures. Section IV presents the experimental results and provides a discussion on the comparative performance of the models. Finally, Section V concludes the paper and outlines potential directions for future research.
Automated histopathology image classification has seen substantial improvements with the advent of deep learning techniques, particularly with convolutional neural networks (CNNs). However, existing methods, such as using generic CNNs like VGGNet and ResNet, may fail to capture critical features from histopathology images due to their complex structure and variability in magnification scales. EfficientNet models have demonstrated strong performance in image classification tasks due to their scalability and efficiency. To further enhance feature extraction, attention mechanisms such as Convolutional Block Attention Module (CBAM), Self-Attention, and Deformable Attention have been proposed.
In this study, we explore Hybrid EfficientNet models that incorporate these advanced attention mechanisms to address the challenges posed by histopathological image variability. The models are evaluated across multiple magnification scales (40X, 100X, 200X, 400X) using the BreakHis dataset, which is a widely recognized dataset for breast cancer histopathology image classification.
II Literature Review
In recent years, extensive research has been conducted on breast cancer classification using deep learning, particularly convolutional neural networks (CNNs). Several studies have explored the use of hybrid models and attention mechanisms to improve the performance of CNNs in histopathology image classification.
Chuang Zhu et al. proposed a hybrid model that assembles multiple compact CNNs for breast cancer classification. This architecture includes a squeeze excitation pruning block and local and global branches, which help reduce channel redundancy and provide a strong representation of the input images. Experiments were performed using the BreakHis and BACH datasets, and the multi-model assembling strategy achieved comparable results with state-of-the-art models Another approach by Togacar et al. developed a residual architecture called BreastNet, which includes attention modules for identifying key regions in histopathology images. The hypercolumn technique was also employed to enhance feature representation. The BreastNet model achieved an accuracy of 98.80% using dense pooling and residual blocks.
Similarly, EM Nejad et al. introduced a single-layer CNN to extract salient features from histopathology images from the BreakHis dataset. Their model achieved a recognition rate of 77.5% . AA Nahid et al. combined Long Short-Term Memory (LSTM) networks with CNNs for breast cancer classification, using Support Vector Machines (SVM) and softmax for final decision-making. Their model achieved 91.00% accuracy.
Yan et al. proposed a model that integrates recurrent neural networks (RNN) with CNNs for histopathology image analysis. Using the BreakHis dataset, they achieved an accuracy of 91.3% for the classification of four cancer classes. Xie et al. also explored deep convolutional neural networks (DCNN) for binary and multi-class breast cancer classification. Their approach used InceptionResNetV2 to extract features, which were further compressed using an autoencoder. This model achieved improved clustering results.
Several studies have leveraged transfer learning to mitigate the data scarcity issue in medical imaging. Sana Ullah et al. developed a model that extracts features from pre-trained networks like VGGNet, GoogleNet, and ResNet. These features are then passed through fully connected layers for binary classification, achieving 97.525% accuracy. Ahmad et al. applied transfer learning using AlexNet, ResNet, and GoogleNet for breast cancer classification, obtaining a maximum accuracy of 85% with ResNet . In another study, Vesal et al. applied transfer learning to classify breast cancer into four sub-classes using the BACH dataset. They utilized ResNet50 and InceptionV3, achieving an accuracy of 97.50% .
Despite these advances, many of the existing methods fail to fully capture the complex structure of breast histopathology images, particularly at different magnification scales. This has led to the exploration of attention mechanisms, which improve the ability of models to focus on the most relevant regions of the image. Woo et al. proposed the Convolutional Block Attention Module (CBAM), which applies attention along both the channel and spatial dimensions to enhance feature extraction . By focusing on critical areas, CBAM has demonstrated significant performance improvements in medical image classification tasks.
In this context, our work aims to further enhance breast cancer classification by integrating attention mechanisms, such as CBAM, Self-Attention, and Deformable Attention, with the EfficientNet architecture. EfficientNet has demonstrated state-of-the-art performance in image classification tasks due to its compound scaling strategy, which optimally balances model depth, width, and resolution. By combining EfficientNet with attention mechanisms, our Hybrid EfficientNet models are designed to focus on both local and global features in histopathology images, leading to improved classification performance.
In this study, we evaluate several Hybrid EfficientNet models integrated with attention mechanisms on the BreakHis dataset, which contains histopathology images at multiple magnification levels. The inclusion of CBAM, Self-Attention, and Deformable Attention in these models enables them to dynamically adjust their focus on key areas of interest in the images, resulting in more accurate and reliable breast cancer classification.
Our proposed approach builds upon existing research by incorporating advanced attention mechanisms and leveraging the strengths of EfficientNet, making it more robust in handling the variability of histopathology images. Through comprehensive experimentation and comparison with other state-of-the-art methods, we aim to demonstrate that our Hybrid EfficientNet models significantly improve breast cancer classification accuracy, particularly at higher magnification scales.
III Methodology
This section outlines the detailed methodology employed in this study for classifying breast cancer histopathology images using Hybrid EfficientNet models integrated with advanced attention mechanisms. Our focus is on enhancing model performance through a robust preprocessing pipeline, modality-specific transfer learning, and a well-defined training strategy.
III-A Hybrid EfficientNet Model Architecture
The backbone of our proposed model consists of the EfficientNet architecture, specifically EfficientNet-B4 and EfficientNet-B8. These models leverage compound scaling to optimize accuracy and efficiency, achieving state-of-the-art results in various image classification tasks. The architecture utilizes a combination of depth, width, and resolution scaling to adapt to the complexities of histopathological images, which often exhibit considerable variability in cellular structures and staining quality.
To enhance feature extraction capabilities, we integrate the following advanced attention mechanisms:
  • Convolutional Block Attention Module (CBAM): The CBAM module consists of two sequential attention mechanisms—channel attention and spatial attention. The channel attention mechanism allows the model to focus on informative features by leveraging global information from feature maps, while spatial attention helps the model concentrate on salient regions within the feature maps. This two-step process is crucial for enhancing the network’s focus on pertinent structures in histopathological images.
  • Self-Attention: The Self-Attention mechanism captures long-range dependencies across the input image, enabling the model to aggregate contextual information from various parts of the image. This is particularly important for histopathology images, where cancerous patterns may be dispersed throughout the tissue.
  • Deformable Attention: This mechanism adaptively adjusts the receptive fields in response to the input features, allowing the model to better capture complex shapes and variations common in histopathological images. The deformable attention layers are instrumental in adapting to the irregularities of tumor shapes and sizes, thus improving classification accuracy.
The integrated architecture, depicted in Fig. 1, begins with the EfficientNet backbone, followed by attention modules that refine the feature representation before passing through fully connected classification layers.
III-B CBAM-EfficientNet Architecture
In this section, we detail the proposed modality-specific CBAM-EfficientNet architecture designed for the classification of H and E stained breast histopathology images. This architecture enhances the traditional EfficientNet framework by integrating the Convolutional Block Attention Module (CBAM), allowing for improved focus on relevant features within the input images.
Figure 1: CBAM
III-B1 Channel Attention Module
The Channel Attention Module (CAM) is pivotal in our architecture for identifying significant features across the channel dimension. This module utilizes the interdependencies among channels, emphasizing crucial information while suppressing less relevant features. The channel attention process follows these steps:
1. Pooling Operations: The input feature map $F$ is first subjected to average pooling and maximum pooling to capture the global contextual information:
$F_{avg}=\text{AvgPool}(F)\quad\text{and}\quad F_{max}=\text{MaxPool}(F)$
2. Feature Aggregation: The pooled feature maps are then fed into a multi-layer perceptron (MLP) to generate channel attention scores, $M_{c}(F)$, which are defined as:
$M_{c}(F)=\sigma\left(\text{MLP}(F_{avg})+\text{MLP}(F_{max})\right)$
where $\sigma$ denotes the sigmoid activation function.
3. Weighting: The resulting channel attention scores are used to weight the original feature map $F$ to obtain the refined feature representation $F^{\prime}_{c}$:
$F^{\prime}_{c}=M_{c}(F)\odot F$
Here, $\odot$ represents the element-wise multiplication, effectively scaling the original features based on their importance.
III-B2 Spatial Attention Module
Following the channel attention mechanism, the Spatial Attention Module (SAM) refines the feature maps by focusing on the spatial locations of the features. This module enhances the model’s capability to discern where important features are located in the input image.
1. Feature Map Aggregation: The spatial attention mechanism aggregates the features across the channel dimension by applying both average and maximum pooling, resulting in two feature maps:
$F_{avg\_c}=\text{AvgPool}(F^{\prime})\quad\text{and}\quad F_{max\_c}=\text{MaxPool}(F^{\prime})$
2. Spatial Attention Calculation: These aggregated maps are then concatenated and passed through a convolutional layer to produce the spatial attention map $M_{s}(F^{\prime})$:
$M_{s}(F^{\prime})=\sigma\left(f_{7\times 7}([F_{avg\_c};F_{max\_c}])\right)$
where $f_{7\times 7}$ denotes a convolution operation with a $7\times 7$ kernel.
3. **Feature Refinement**: The refined feature map $F^{\prime}_{s}$ is obtained by applying the spatial attention map to the already refined channel-wise features:
$F^{\prime}_{s}=M_{s}(F^{\prime})\odot F^{\prime}$
Figure 2: Spatial and Channel Attention
III-B3 EfficientNet Backbone
The backbone of our model utilizes the EfficientNet architecture, known for its efficiency in parameter utilization and computational performance. Each EfficientNet variant is constructed using a combination of depthwise separable convolutions and squeeze-and-excitation blocks, allowing for enhanced feature extraction while maintaining low computational costs.
We integrate the CBAM modules at various points in the EfficientNet architecture to enhance the focus on the most relevant features. The output from the final layer is then passed through a Global Average Pooling (GAP) layer, which reduces the dimensionality while preserving important spatial information:
$F_{GAP}=\text{GAP}(F^{\prime}_{s})$
Finally, the output from the GAP layer is fed into fully connected layers for classification tasks, ensuring that the model captures both local and global features effectively.
This combined approach of CBAM and EfficientNet leverages both channel and spatial attention mechanisms, providing a robust framework for the classification of breast histopathology images while maintaining computational efficiency.
III-C Datasets
For our experiments, we utilized the BreakHis dataset, which consists of 9,109 breast histopathology images sourced from 82 patients. The dataset contains images at four magnification levels: 40X, 100X, 200X, and 400X, categorized into benign and malignant classes. To enhance model generalizability, we also employed additional datasets such as ICIAR 2018 and the PCam dataset for pre-training. The ICIAR 2018 dataset includes a variety of histopathology images labeled for benign and malignant conditions, while the PCam dataset provides a substantial number of whole-slide images with binary labels indicating metastatic tissues.
III-D Data Preprocessing
To address the challenges posed by histopathology images, we implemented a comprehensive hybrid preprocessing pipeline. The following steps were undertaken:
  • Zero Padding: We applied zero padding to handle edge pixels effectively, ensuring that convolutional operations do not lose critical information at the borders of images. This technique adds additional rows and columns around the image, extending beyond its boundary.
  • Median Filtering: To reduce additive noise while preserving the structural integrity of the tissue, we employed a median filter. This non-linear filter replaces each pixel value with the median of its neighborhood values, effectively removing noise without blurring edges, which are vital for accurate feature extraction.
  • Contrast Limited Adaptive Histogram Equalization (CLAHE): Given the uneven staining in histopathology images, CLAHE was utilized to enhance local contrast and improve weak boundary detection. By dividing the image into small tiles and applying histogram equalization to each, we improve the visibility of critical features within the images.
  • Normalization: Image pixel values were normalized to the range [0, 1] to enhance model convergence during training. This ensures uniformity in the input data distribution, which is crucial for the performance of deep learning models.
  • Data Augmentation: To mitigate the challenges of limited sample sizes, extensive data augmentation was performed on the BreakHis dataset. Techniques such as rotation, flipping, zooming, and brightness adjustment were employed to artificially increase the dataset size and introduce variability, thereby reducing the risk of overfitting.
III-E Training and Hyperparameter Tuning
The training of the models was conducted using PyTorch on an NVIDIA RTX 4060 GPU, enabling efficient computation. We utilized an initial learning rate of 0.001 and employed both the Adam and SGD optimizers to determine which provided better convergence and generalization.
Key hyperparameters were carefully tuned:
  • Dense Layers: Experiments were conducted with two and three dense layers to identify the optimal configuration for the final classifier.
  • Activation Functions: Both ReLU and tanh activation functions were tested at the last classification layer to evaluate their impact on the model’s performance.
  • Kernel Initializers: We experimented with different kernel initializers, including the normal initializer and He initialization, to determine their effects on the convergence rate during training.
  • Dropout Regularization: Dropout was incorporated in the fully connected layers to prevent overfitting, particularly given the limited availability of labeled data in medical imaging tasks.
  • Batch Size and Epochs: We tested various batch sizes (16, 32, and 64) and trained the models for a maximum of 100 epochs, applying early stopping based on validation loss to prevent overfitting.
III-F Modality-Specific Transfer Learning Strategy
Given the inherent limitations of training deep learning models with small datasets, we employed a modality-specific transfer learning approach. This strategy leverages pre-trained models on large, similar datasets to enhance feature extraction capabilities for our specific task.
We fine-tuned the EfficientNet models, initially pre-trained on diverse cancer datasets (PCam, ICIAR 2018, and others) rather than the ImageNet dataset, to reduce the domain gap inherent in medical imaging tasks. By tailoring the transfer learning process to domain-specific data, we aimed to improve model robustness and classification accuracy on the BreakHis dataset.
III-G Evaluation Metrics
The performance of our proposed models was assessed using a suite of evaluation metrics: accuracy, precision, recall, and F1-score. These metrics are crucial in medical applications where both false negatives and false positives can have severe implications for patient care. Moreover, the performance was evaluated across different magnification levels (40X, 100X, 200X, and 400X) to analyze how well the models generalize to various scales of tissue detail.
By employing this detailed methodology, we aim to advance the field of breast cancer histopathology classification and contribute to the development of more reliable and efficient diagnostic tools.
TABLE I: Performance of Various Models Across Magnification Scales
Model Scale Accuracy (%) F1 Score (%) Precision Recall
Hybrid EfficientNet WideSEB4 + CBAM 40X 96.237 95.781 0.963 0.956
100X 97.115 96.863 0.971 0.970
200X 97.865 97.672 0.979 0.979
400X 98.423 97.951 0.985 0.983
Hybrid EfficientNet B8 + CBAM 40X 95.564 95.323 0.957 0.951
100X 97.391 96.911 0.973 0.975
200X 98.091 97.723 0.980 0.980
400X 98.303 97.834 0.988 0.981
Hybrid EfficientNet WideSEB4 + Self-Attention 40X 94.506 94.234 0.947 0.942
100X 96.804 96.453 0.968 0.966
200X 97.305 96.983 0.973 0.972
400X 97.856 97.582 0.979 0.978
Hybrid EfficientNet B8 + Self-Attention 40X 93.806 93.531 0.939 0.936
100X 96.402 96.023 0.964 0.962
200X 97.108 96.645 0.971 0.970
400X 97.657 97.321 0.978 0.975
EfficientNet-B4-WideSE without CBAM 40X 84.754 84.476 0.843 0.848
100X 85.852 85.539 0.856 0.858
200X 86.902 86.774 0.868 0.869
400X 87.503 87.434 0.874 0.875
EfficientNet-B8 without CBAM 40X 84.603 84.153 0.841 0.846
100X 85.901 86.225 0.865 0.863
200X 87.204 87.411 0.875 0.876
400X 87.901 87.673 0.878 0.879
EfficientNet-B4 with Deformable Attention 40X 95.204 95.124 0.952 0.954
100X 95.901 95.916 0.962 0.960
200X 96.301 96.268 0.965 0.963
400X 96.604 96.534 0.968 0.966
EfficientNet-B8 with Deformable Attention 40X 95.404 95.372 0.959 0.954
100X 96.002 96.115 0.963 0.961
200X 96.402 96.442 0.967 0.964
400X 96.705 96.653 0.969 0.968
IV Experimental Results and Discussions
IV-A Experimental Platform
The experiments were conducted using PyTorch as the deep learning framework backend. The models were trained and tested on an NVIDIA RTX 4060 GPU, providing substantial computational power for the deep learning tasks involved. The system was equipped with 16GB of RAM to efficiently handle the large dataset and facilitate faster processing.
IV-B Hyperparameter Optimization
Hyperparameters are critical variables that govern the training process of convolutional neural networks (CNNs). While the networks learn the relationships between inputs and outputs, optimizing hyperparameters is essential to enhance model performance. We performed extensive experiments to identify the optimal hyperparameters for our final training.
We evaluated two different optimizers: Adam (Adaptive Moment Estimation) and SGD (Stochastic Gradient Descent), with three distinct learning rates ($0.01$, $0.001$, and $0.0001$). The SoftMax classifier was employed for all classification experiments, providing probability scores for each class label.
Binary cross-entropy was used as the loss function due to the binary nature of our classification problem, represented mathematically as follows:
[…]
The fully connected layer incorporated a ReLU activation function and consisted of 256 hidden neurons, followed by a dropout layer with a probability of $0.4$ to mitigate overfitting.
IV-C Results
We employed a modality-specific transfer learning strategy for the classification task and compared it with various state-of-the-art CNN architectures, as shown in Table II. Multiple experiments were conducted to evaluate the binary classification of breast cancer histopathology images across various magnification factors (40X, 100X, 200X, and 400X).
To demonstrate the efficacy of our proposed approach, we compared the performance of several models and provided a detailed comparative analysis. Initially, we conducted experiments on two different models without transfer learning, as presented in Table 1. Subsequently, we evaluated state-of-the-art models pre-trained on ImageNet, as summarized. The performance of our proposed modality-specific strategy, incorporating transfer learning.
As shown in Table 2, our proposed Hybrid EfficientNet WideSEB4 + CBAM significantly outperformed various established models, including DenseNet201, and Deep CNNs. Notably, at higher magnification levels, our hybrid model achieved an accuracy of 98.42% at 400X magnification.
IV-C1 Comparison of Hybrid EfficientNet vs VGG16 + VGG19
The Hybrid EfficientNet WideSEB4 + CBAM demonstrated significant performance improvements over the VGG16 + VGG19 combination across all magnifications. While the VGG-based model performed well in terms of accuracy, its computational requirements were substantially higher. In our experiments, the VGG16 + VGG19 model required the use of a Tesla K80 GPU, which provides substantial computational resources and a larger memory capacity of 24GB VRAM, essential for handling the large parameter count of the combined models.
In contrast, the Hybrid EfficientNet WideSEB4 + CBAM was trained on an NVIDIA RTX 4060 GPU, which has only 8GB of VRAM—considerably less than the Tesla K80. Despite the significantly lower memory, the EfficientNet-based architecture demonstrated exceptional computational efficiency, primarily due to its utilization of depthwise separable convolutions, squeeze-and-excitation (SE) blocks, and the Convolutional Block Attention Module (CBAM). With a lower parameter count and memory requirements, EfficientNet maintained superior accuracy, particularly at higher magnifications.
TABLE II: Comparison of Accuracy Across Models and Magnification Factors
Method Model MG Factor Accuracy (%)
Erfankhan Hamed, et al LBP 40X 88.30
100X 88.30
200X 87.10
400X 83.40
Daniel Lichtblau, et al AlexNet 40X 81.61
100X 84.47
200X 86.67
400X 83.15
Yun Gu, et al DCMM 40X 95.62
100X 95.03
200X 97.04
400X 96.31
Nahid, et al Deep CNN 40X 90.00
100X 91.00
200X 91.00
400X 90.00
Togacar Mesut, et al Deep CNN 40X 97.99
100X 97.84
200X 98.51
400X 95.88
Yan Hao, et al DenseNet 201 40X 96.75
100X 95.21
200X 96.57
400X 93.15
Pin Wang, et al FE-BkCapsNet 40X 92.71
100X 94.52
200X 94.03
400X 93.54
VGG16 + VGG19 Hameed et al 40X 94.44
100X 97.61
200X 98.70
400X 98.96
Proposed Hybrid 2024 40X 97.11
(EfficientNet WideSEB4 + CBAM) 100X 98.04
200X 98.25
400X 98.42
IV-C2 Parameter Efficiency and Memory Consumption
EfficientNet leverages a compound scaling method to optimize model depth, width, and input resolution, allowing it to achieve high accuracy with fewer parameters compared to VGG models. Specifically:
  • The combined VGG16 + VGG19 model consists of approximately 40 million parameters, necessitating substantial memory resources and extended training times.
  • In contrast, the Hybrid EfficientNet WideSEB4 + CBAM model, even with the inclusion of the CBAM module, contains fewer than 30 million parameters, leading to significantly lower memory consumption and faster computation. This reduction in parameter count allowed for the training of the EfficientNet model on an RTX 4060 GPU, despite its more limited memory capacity of 8GB. Furthermore, EfficientNet’s superior optimization led to faster convergence and improved accuracy, making it a more feasible choice for large-scale applications where computational resources are constrained.
IV-C3 Best Model Performance
Hybrid EfficientNet WideSEB4 + CBAM achieved the 2nd highest accuracy of 98.42% at 400X magnification, highlighting the model’s capability in extracting detailed and relevant features from breast cancer histopathology images. In comparison with other state-of-the-art models, the proposed architecture not only surpassed accuracy benchmarks but also offered substantial advantages in terms of computational efficiency and memory requirements, underscoring its potential for deployment in real-time clinical settings.
V Conclusion
This study presents a comprehensive approach for breast cancer histopathology classification using Hybrid EfficientNet models integrated with CBAM, Self-Attention, and Deformable Attention mechanisms. The proposed model achieved superior accuracy and computational efficiency across multiple magnification levels, with peak performance at 400X magnification. By leveraging attention mechanisms to focus on critical regions, our approach effectively addresses the complexity of histopathological images. These results demonstrate the potential of our model for integration into real-time diagnostic workflows, offering enhanced diagnostic accuracy and efficiency. Future work could explore further model adaptations to other histopathology datasets and expand clinical applicability.

### v3

@@
-EfficientNet with Hybrid Attention Mechanisms for Enhanced Breast Histopathology Classification: A Comprehensive Approach
+Enhanced Histopathology Image Feature Extraction using EfficientNet with Dual Attention Mechanisms and CLAHE Preprocessing
 Abstract
-Breast cancer histopathology image classification is crucial for early cancer detection, offering the potential to reduce mortality rates through timely diagnosis.
-This paper introduces a novel approach integrating Hybrid EfficientNet models with advanced attention mechanisms, including Convolutional Block Attention Module (CBAM), Self-Attention, and Deformable Attention, to enhance feature extraction and focus on critical image regions.
-We evaluate the performance of our models across multiple magnification scales using publicly available histopathological datasets.
-Our method achieves significant improvements, with accuracy reaching 98.42% at 400X magnification, surpassing several state-of-the-art models, including VGG and ResNet architectures.
-The results are validated using metrics such as accuracy, F1-score, precision, and recall, demonstrating the clinical potential of our model in improving diagnostic accuracy.
-Furthermore, the proposed method shows increased computational efficiency, making it suitable for integration into real-time diagnostic workflows.
+Breast cancer histopathology image classification is critical for early detection and improved patient outcomes.
+This study introduces a novel approach leveraging Hybrid EfficientNet models enhanced with advanced attention mechanisms—CBAM, Self-Attention, and Deformable Attention—to improve feature extraction and focus on relevant tissue regions.
+The proposed models were evaluated on the BreakHis dataset across multiple magnification scales (40X, 100X, 200X, and 400X).
+Among them, the EfficientNetV2-XL with CBAM achieved outstanding performance, reaching a peak accuracy of 98.96% and an F1-score of 98.31% at 400X magnification, outperforming state-of-the-art methods.
+By integrating Contrast Limited Adaptive Histogram Equalization (CLAHE) for preprocessing and optimizing computational efficiency, this method demonstrates its suitability for real-time clinical deployment.
+The results underscore the potential of attention-enhanced scalable architectures in advancing diagnostic precision for breast cancer detection.
 Index Terms:
@@
 EfficientNet has demonstrated state-of-the-art performance in image classification tasks due to its compound scaling strategy, which optimally balances model depth, width, and resolution.
-By combining EfficientNet with attention mechanisms, our Hybrid EfficientNet models are designed to focus on both local and global features in histopathology images, leading to improved classification performance.
-In this study, we evaluate several Hybrid EfficientNet models integrated with attention mechanisms on the BreakHis dataset, which contains histopathology images at multiple magnification levels.
-The inclusion of CBAM, Self-Attention, and Deformable Attention in these models enables them to dynamically adjust their focus on key areas of interest in the images, resulting in more accurate and reliable breast cancer classification.
 Our proposed approach builds upon existing research by incorporating advanced attention mechanisms and leveraging the strengths of EfficientNet, making it more robust in handling the variability of histopathology images.
@@
 III-A Hybrid EfficientNet Model Architecture
-The backbone of our proposed model consists of the EfficientNet architecture, specifically EfficientNet-B4 and EfficientNet-B8.
+The backbone of our proposed model consists of the EfficientNet architecture, specifically EfficientNetV2-XL and EfficientNetV2-L.
 These models leverage compound scaling to optimize accuracy and efficiency, achieving state-of-the-art results in various image classification tasks.
@@
 This combined approach of CBAM and EfficientNet leverages both channel and spatial attention mechanisms, providing a robust framework for the classification of breast histopathology images while maintaining computational efficiency.
+Figure 3: Visualization of Attention Mechanisms in Histopathological Images
 III-C Datasets
@@
 By tailoring the transfer learning process to domain-specific data, we aimed to improve model robustness and classification accuracy on the BreakHis dataset.
+TABLE I: Performance of Base Models Across Magnification Scales
+Model Scale Accuracy (%) F1 Score (%)
+EfficientNetV2-XL (Base) 40X 86.78 84.23
+100X 87.89 85.34
+200X 88.23 86.67
+400X 89.12 87.56
+EfficientNetV2-L (Base) 40X 85.67 83.13
+100X 86.89 84.23
+200X 87.13 85.56
+400X 87.89 86.23
 III-G Evaluation Metrics
@@
 Moreover, the performance was evaluated across different magnification levels (40X, 100X, 200X, and 400X) to analyze how well the models generalize to various scales of tissue detail.
-By employing this detailed methodology, we aim to advance the field of breast cancer histopathology classification and contribute to the development of more reliable and efficient diagnostic tools.
-TABLE I: Performance of Various Models Across Magnification Scales
-Model Scale Accuracy (%) F1 Score (%) Precision Recall
-Hybrid EfficientNet WideSEB4 + CBAM 40X 96.237 95.781 0.963 0.956
-100X 97.115 96.863 0.971 0.970
-200X 97.865 97.672 0.979 0.979
-400X 98.423 97.951 0.985 0.983
-Hybrid EfficientNet B8 + CBAM 40X 95.564 95.323 0.957 0.951
-100X 97.391 96.911 0.973 0.975
-200X 98.091 97.723 0.980 0.980
-400X 98.303 97.834 0.988 0.981
-Hybrid EfficientNet WideSEB4 + Self-Attention 40X 94.506 94.234 0.947 0.942
-100X 96.804 96.453 0.968 0.966
-200X 97.305 96.983 0.973 0.972
-400X 97.856 97.582 0.979 0.978
-Hybrid EfficientNet B8 + Self-Attention 40X 93.806 93.531 0.939 0.936
-100X 96.402 96.023 0.964 0.962
-200X 97.108 96.645 0.971 0.970
-400X 97.657 97.321 0.978 0.975
-EfficientNet-B4-WideSE without CBAM 40X 84.754 84.476 0.843 0.848
-100X 85.852 85.539 0.856 0.858
-200X 86.902 86.774 0.868 0.869
-400X 87.503 87.434 0.874 0.875
-EfficientNet-B8 without CBAM 40X 84.603 84.153 0.841 0.846
-100X 85.901 86.225 0.865 0.863
-200X 87.204 87.411 0.875 0.876
-400X 87.901 87.673 0.878 0.879
-EfficientNet-B4 with Deformable Attention 40X 95.204 95.124 0.952 0.954
-100X 95.901 95.916 0.962 0.960
-200X 96.301 96.268 0.965 0.963
-400X 96.604 96.534 0.968 0.966
-EfficientNet-B8 with Deformable Attention 40X 95.404 95.372 0.959 0.954
-100X 96.002 96.115 0.963 0.961
-200X 96.402 96.442 0.967 0.964
-400X 96.705 96.653 0.969 0.968
+Figure 4: ROC-AUC Curve of EfficientNetV2-XL with CBAM
+TABLE II: Performance of Various Models Across Magnification Scales
+Model Scale Accuracy (%) F1 Score (%)
+EfficientNetV2-XL + CBAM 40X 95.02 94.13
+100X 96.85 95.69
+200X 97.45 96.96
+400X 98.96 98.31
+EfficientNetV2-L + CBAM 40X 94.89 97.78
+100X 96.13 97.99
+200X 97.85 98.03
+400X 99.16 98.15
+EfficientNetV2-XL + Self-Attention 40X 96.56 92.02
+100X 97.35 93.29
+200X 97.13 94.16
+400X 98.56 95.43
+EfficientNetV2-L + Self-Attention 40X 96.23 91.89
+100X 97.13 92.78
+200X 97.89 93.93
+400X 98.45 94.89
+EfficientNetV2-XL + Deformable Attention 40X 96.35 95.89
+100X 96.89 96.35
+200X 97.13 96.56
+400X 97.56 96.93
+EfficientNetV2-L + Deformable Attention 40X 96.13 94.17
+100X 96.78 94.23
+200X 96.89 95.35
+400X 97.45 96.49
 IV Experimental Results and Discussions
@@
 IV-C Results
-We employed a modality-specific transfer learning strategy for the classification task and compared it with various state-of-the-art CNN architectures, as shown in Table II.
-Multiple experiments were conducted to evaluate the binary classification of breast cancer histopathology images across various magnification factors (40X, 100X, 200X, and 400X).
-To demonstrate the efficacy of our proposed approach, we compared the performance of several models and provided a detailed comparative analysis.
+We employed a modality-specific transfer learning strategy for the classification task and compared it with various state-of-the-art CNN architectures, as shown in Table III.
 Initially, we conducted experiments on two different models without transfer learning, as presented in Table 3.
@@
 The performance of our proposed modality-specific strategy, incorporating transfer learning.
-As shown in Table 2, our proposed Hybrid EfficientNet WideSEB4 + CBAM significantly outperformed various established models, including DenseNet201, and Deep CNNs.
-Notably, at higher magnification levels, our hybrid model achieved an accuracy of 98.42% at 400X magnification.
-IV-C1 Comparison of Hybrid EfficientNet vs VGG16 + VGG19
-The Hybrid EfficientNet WideSEB4 + CBAM demonstrated significant performance improvements over the VGG16 + VGG19 combination across all magnifications.
-While the VGG-based model performed well in terms of accuracy, its computational requirements were substantially higher.
-In our experiments, the VGG16 + VGG19 model required the use of a Tesla K80 GPU, which provides substantial computational resources and a larger memory capacity of 24GB VRAM, essential for handling the large parameter count of the combined models.
-In contrast, the Hybrid EfficientNet WideSEB4 + CBAM was trained on an NVIDIA RTX 4060 GPU, which has only 8GB of VRAM—considerably less than the Tesla K80.
-Despite the significantly lower memory, the EfficientNet-based architecture demonstrated exceptional computational efficiency, primarily due to its utilization of depthwise separable convolutions, squeeze-and-excitation (SE) blocks, and the Convolutional Block Attention Module (CBAM).
-With a lower parameter count and memory requirements, EfficientNet maintained superior accuracy, particularly at higher magnifications.
-TABLE II: Comparison of Accuracy Across Models and Magnification Factors
+As shown in Table 2, our proposed EfficientNetV2-XL + CBAM significantly outperformed various established models, including DenseNet201 and Deep CNNs.
+Notably, at higher magnification levels, this model achieved an outstanding accuracy of 98.96% and an F1-score of 98.31% at 400X magnification, highlighting its ability to accurately classify breast cancer histopathology images
+The EfficientNetV2-XL + CBAM demonstrated substantial performance improvements over the VGG16 + VGG19 combination across all magnifications.
+While the VGG-based models delivered competitive accuracy, their computational requirements were significantly higher.
+For instance, the VGG16 + VGG19 models required a Tesla K80 GPU with 24GB of VRAM to accommodate their large parameter count and high computational demand.
+In contrast, the EfficientNetV2-XL + CBAM was trained on a more modest NVIDIA RTX 4060 GPU with just 8GB of VRAM.
+Despite the lower memory availability, the EfficientNetV2-XL architecture demonstrated exceptional computational efficiency, leveraging depthwise separable convolutions, squeeze-and-excitation (SE) blocks, and the Convolutional Block Attention Module (CBAM).
+This efficiency allowed the model to maintain superior accuracy and scalability, particularly at higher magnifications, while requiring significantly fewer computational resources.
+TABLE III: Comparison of Accuracy Across Models and Magnification Factors
 Method Model MG Factor Accuracy (%)
@@
 Proposed Hybrid 2024 40X 97.11
-(EfficientNet WideSEB4 + CBAM) 100X 98.04
+(EfficientNetV2-XL + CBAM) 100X 98.04
 200X 98.25
-400X 98.42
-IV-C2 Parameter Efficiency and Memory Consumption
-EfficientNet leverages a compound scaling method to optimize model depth, width, and input resolution, allowing it to achieve high accuracy with fewer parameters compared to VGG models. Specifically:
-  • The combined VGG16 + VGG19 model consists of approximately 40 million parameters, necessitating substantial memory resources and extended training times.
-  • In contrast, the Hybrid EfficientNet WideSEB4 + CBAM model, even with the inclusion of the CBAM module, contains fewer than 30 million parameters, leading to significantly lower memory consumption and faster computation.
-This reduction in parameter count allowed for the training of the EfficientNet model on an RTX 4060 GPU, despite its more limited memory capacity of 8GB.
-Furthermore, EfficientNet’s superior optimization led to faster convergence and improved accuracy, making it a more feasible choice for large-scale applications where computational resources are constrained.
-IV-C3 Best Model Performance
-Hybrid EfficientNet WideSEB4 + CBAM achieved the 2nd highest accuracy of 98.42% at 400X magnification, highlighting the model’s capability in extracting detailed and relevant features from breast cancer histopathology images.
-In comparison with other state-of-the-art models, the proposed architecture not only surpassed accuracy benchmarks but also offered substantial advantages in terms of computational efficiency and memory requirements, underscoring its potential for deployment in real-time clinical settings.
+400X 98.96
 V Conclusion
-This study presents a comprehensive approach for breast cancer histopathology classification using Hybrid EfficientNet models integrated with CBAM, Self-Attention, and Deformable Attention mechanisms.
-The proposed model achieved superior accuracy and computational efficiency across multiple magnification levels, with peak performance at 400X magnification.
-By leveraging attention mechanisms to focus on critical regions, our approach effectively addresses the complexity of histopathological images.
-These results demonstrate the potential of our model for integration into real-time diagnostic workflows, offering enhanced diagnostic accuracy and efficiency.
-Future work could explore further model adaptations to other histopathology datasets and expand clinical applicability.
+In this study, we propose a Hybrid EfficientNetV2 models integrated with CBAM, Self-Attention, and Deformable Attention mechanisms for breast cancer histopathology classification.
+The EfficientNetV2-XL model, with approximately 208 million parameters and the inclusion of the CBAM module, offers a more efficient and accurate solution compared to previous state-of-the-art models like VGG16 + VGG19, which have around 282 million parameters.
+This reduction in parameters enables faster training and reduced memory consumption, allowing the model to be trained on an RTX 4060 GPU with 8GB of memory.
+The Hybrid EfficientNetV2-XL + CBAM and previous SOTA VGG16 + VGG19 model achieved superior accuracy, however EfficientNetv2-XL has F1-Score of 98.31 and Hybrid VGG16 + VG199 has 97.95% ,therefore EfficientNetV2-XL + CBAM outperforms other models in both accuracy and computational efficiency.
+By leveraging attention mechanisms, the model effectively focuses on critical regions of histopathological images, making it well-suited for real-time clinical diagnostics.
+These results demonstrate that the proposed model surpasses the performance of previous state-of-the-art models, highlighting its potential for real-time clinical deployment.
+Future work may adapt the model to other datasets and further expand its clinical applicability.

### v4

@@
-Enhanced Histopathology Image Feature Extraction using EfficientNet with Dual Attention Mechanisms and CLAHE Preprocessing
+CBAM-EfficientNetV2 for Histopathology Image Classification using Transfer Learning and Dual Attention Mechanisms
 Abstract
 Breast cancer histopathology image classification is critical for early detection and improved patient outcomes.
-This study introduces a novel approach leveraging Hybrid EfficientNet models enhanced with advanced attention mechanisms—CBAM, Self-Attention, and Deformable Attention—to improve feature extraction and focus on relevant tissue regions.
+This study introduces a novel approach leveraging EfficientNetV2 models, to improve feature extraction and focus on relevant tissue regions.
 The proposed models were evaluated on the BreakHis dataset across multiple magnification scales (40X, 100X, 200X, and 400X).
@@
 Index Terms:
-Transfer Learning, Modality Specific, EfficientNet, CBAM, Self-Attention, Deformable Attention, Histopathology, Image Classification, Deep Learning, Breast Cancer.
+Transfer Learning, Modality Specific, EfficientNet, CBAM, Histopathology, Image Classification, Deep Learning, Breast Cancer.
 I Introduction
-Breast cancer remains one of the most prevalent and deadly forms of cancer, especially affecting women worldwide.
-According to global cancer statistics, breast cancer accounts for more than 2.3 million new cases each year, contributing to a significant number of cancer-related deaths [3].
-Early detection plays a critical role in reducing the mortality rate associated with this disease.
-The standard diagnostic procedure for detecting breast cancer involves histopathological analysis, where biopsy tissue samples are examined under a microscope.
-This process is time-consuming, labor-intensive, and prone to subjective interpretation errors, even among trained pathologists [1].
-Pathologists rely on morphological abnormalities in cell structures, particularly the nuclei, to distinguish between benign and malignant cells [2].
-Benign tumors typically exhibit well-organized cellular structures and rarely invade surrounding tissues or metastasize to other parts of the body.
-In contrast, malignant tumors display uncontrolled growth, irregular nuclei, and the potential to spread, making them life-threatening.
-Early identification of these malignant traits is critical for improving patient prognosis and guiding treatment plans.
-In recent years, computer-aided diagnosis (CAD) systems have been developed to assist pathologists by automating the analysis of histopathology images.
-Deep learning (DL) models, particularly Convolutional Neural Networks (CNNs), have shown great promise in this domain due to their ability to automatically learn complex features from medical images [7].
-However, despite their success in natural image classification, traditional CNN architectures, such as VGGNet and ResNet, struggle to capture the intricate features present in histopathological images, primarily due to the variability in magnification scales and the complex structure of the tissue samples.
-To address these challenges, researchers have begun exploring hybrid models that combine CNNs with attention mechanisms.
-These attention mechanisms, including the Convolutional Block Attention Module (CBAM), Self-Attention, and Deformable Attention, have been designed to enhance feature extraction by allowing the network to focus on the most relevant parts of the image.
-CBAM, for example, has shown considerable potential in medical image analysis by directing the network’s attention toward critical regions, leading to improved classification accuracy.
-Similarly, Self-Attention mechanisms, initially introduced in transformer models for Natural Language Processing (NLP), have been adapted to medical imaging tasks, where they capture long-range dependencies across the image.
-Deformable Attention further enhances this by dynamically adjusting the receptive field based on the complexity of the input features, making it particularly effective for identifying abnormalities in histopathology images.
-In this study, we propose a series of Hybrid EfficientNet models integrated with advanced attention mechanisms to improve the performance of breast cancer histopathology image classification.
-EfficientNet, a highly scalable CNN architecture, has been recognized for its ability to achieve state-of-the-art performance with fewer parameters by optimizing the depth, width, and resolution of the model.
-By incorporating attention mechanisms such as CBAM, Self-Attention, and Deformable Attention into EfficientNet, we aim to enhance the network’s focus on key features, thereby improving the classification of benign and malignant breast cancer tissue samples.
-The dataset used in this study is the BreakHis dataset, a publicly available dataset widely used for breast cancer histopathology image classification tasks .
-The dataset consists of 9,109 images at four different magnification levels (40X, 100X, 200X, and 400X), providing a comprehensive set of data to test the proposed models across varying scales of tissue detail.
-The performance of the proposed models is evaluated using key metrics such as accuracy, precision, recall, and F1-score.
-The rest of the paper is structured as follows: In Section II, we review the existing literature on breast cancer histopathology image classification and attention mechanisms.
-Section III describes the methodology, including the dataset, preprocessing techniques, model architecture, and training procedures.
-Section IV presents the experimental results and provides a discussion on the comparative performance of the models.
-Finally, Section V concludes the paper and outlines potential directions for future research.
-Automated histopathology image classification has seen substantial improvements with the advent of deep learning techniques, particularly with convolutional neural networks (CNNs).
-However, existing methods, such as using generic CNNs like VGGNet and ResNet, may fail to capture critical features from histopathology images due to their complex structure and variability in magnification scales.
-EfficientNet models have demonstrated strong performance in image classification tasks due to their scalability and efficiency.
-To further enhance feature extraction, attention mechanisms such as Convolutional Block Attention Module (CBAM), Self-Attention, and Deformable Attention have been proposed.
-In this study, we explore Hybrid EfficientNet models that incorporate these advanced attention mechanisms to address the challenges posed by histopathological image variability.
-The models are evaluated across multiple magnification scales (40X, 100X, 200X, 400X) using the BreakHis dataset, which is a widely recognized dataset for breast cancer histopathology image classification.
+Automating histopathological analysis has become a vital research area in medical imaging due to its potential to significantly reduce the diagnostic workload of pathologists and enhance the consistency and reproducibility of results.
+Histopathology—the microscopic study of tissue abnormalities—is fundamental in diagnosing diseases such as cancer, but manual analysis is labor-intensive, subjective, and prone to variability between observers.
+As a result, computer-aided diagnosis (CAD) systems have gained traction, with the aim of augmenting pathologists’ abilities to efficiently and accurately classify tissue samples [1].
+Despite substantial progress, CAD systems continue to face significant challenges, particularly in handling the variability inherent in histopathological images.
+Differences in magnification levels, staining protocols (e.g., Hematoxylin and Eosin, or H&E staining), and tissue morphology introduce variability that complicates the generalization of standard deep learning models [2].
+Traditional convolutional neural networks (CNNs), while successful in many computer vision tasks, often struggle to generalize across these variations in histopathological data.
+Specifically, CNNs are designed with fixed receptive fields and are limited in their ability to capture fine-grained details that vary across multiple scales [3].
+The heterogeneity of breast cancer histopathology images—marked by different magnification levels and complex cellular structures—poses unique challenges that demand more sophisticated architectures.
+To address this, researchers have turned to hybrid models that integrate attention mechanisms with CNNs, which have demonstrated superior performance in medical imaging tasks by enhancing feature extraction.
+Attention mechanisms such as the Convolutional Block Attention Module (CBAM) have been adapted to improve image classification by enabling the network to focus dynamically on the most informative regions of an image.
+In histopathology, these mechanisms are particularly effective in highlighting regions critical to cancer diagnosis, such as irregular nuclei shapes, abnormal cellular arrangements, and other key tissue structures that may indicate malignancy.
+By applying attention, the network can dynamically recalibrate feature maps along spatial and channel dimensions, improving its ability to capture both local and global image features [4].
+This is especially valuable in distinguishing between benign and malignant tissue, where small but significant differences in morphology must be detected.
+EfficientNet [8], a state-of-the-art CNN architecture known for its compound scaling strategy, is particularly well-suited for medical image classification tasks due to its ability to balance model depth, width, and resolution.
+EfficientNet achieves superior performance with fewer parameters, making it computationally efficient without sacrificing accuracy.
+By integrating attention mechanisms with EfficientNet, we aim to build models that can more effectively classify breast cancer histopathology images while remaining scalable to real-world clinical applications.
+This hybrid approach addresses the need for models that can operate across varying image resolutions and magnification levels, ensuring robustness and high accuracy in diverse settings.
 II Literature Review
-In recent years, extensive research has been conducted on breast cancer classification using deep learning, particularly convolutional neural networks (CNNs).
-Several studies have explored the use of hybrid models and attention mechanisms to improve the performance of CNNs in histopathology image classification.
-Chuang Zhu et al.
-proposed a hybrid model that assembles multiple compact CNNs for breast cancer classification.
-This architecture includes a squeeze excitation pruning block and local and global branches, which help reduce channel redundancy and provide a strong representation of the input images.
-Experiments were performed using the BreakHis and BACH datasets, and the multi-model assembling strategy achieved comparable results with state-of-the-art models Another approach by Togacar et al.
-developed a residual architecture called BreastNet, which includes attention modules for identifying key regions in histopathology images.
-The hypercolumn technique was also employed to enhance feature representation.
-The BreastNet model achieved an accuracy of 98.80% using dense pooling and residual blocks.
-Similarly, EM Nejad et al.
-introduced a single-layer CNN to extract salient features from histopathology images from the BreakHis dataset.
-Their model achieved a recognition rate of 77.5% .
-AA Nahid et al.
-combined Long Short-Term Memory (LSTM) networks with CNNs for breast cancer classification, using Support Vector Machines (SVM) and softmax for final decision-making.
-Their model achieved 91.00% accuracy.
-Yan et al.
-proposed a model that integrates recurrent neural networks (RNN) with CNNs for histopathology image analysis.
-Using the BreakHis dataset, they achieved an accuracy of 91.3% for the classification of four cancer classes.
-Xie et al.
-also explored deep convolutional neural networks (DCNN) for binary and multi-class breast cancer classification.
-Their approach used InceptionResNetV2 to extract features, which were further compressed using an autoencoder.
-This model achieved improved clustering results.
-Several studies have leveraged transfer learning to mitigate the data scarcity issue in medical imaging.
-Sana Ullah et al.
-developed a model that extracts features from pre-trained networks like VGGNet, GoogleNet, and ResNet.
-These features are then passed through fully connected layers for binary classification, achieving 97.525% accuracy.
-Ahmad et al.
-applied transfer learning using AlexNet, ResNet, and GoogleNet for breast cancer classification, obtaining a maximum accuracy of 85% with ResNet .
-In another study, Vesal et al.
-applied transfer learning to classify breast cancer into four sub-classes using the BACH dataset.
-They utilized ResNet50 and InceptionV3, achieving an accuracy of 97.50% .
-Despite these advances, many of the existing methods fail to fully capture the complex structure of breast histopathology images, particularly at different magnification scales.
-This has led to the exploration of attention mechanisms, which improve the ability of models to focus on the most relevant regions of the image.
+Hybrid models have gained significant attention in breast cancer classification due to their ability to combine the strengths of multiple architectures.
+A multi-model assembling strategy that integrates local and global branches, reducing redundancy and enhancing feature representation.
+This approach has achieved high accuracy on standard breast cancer datasets, demonstrating the potential of hybrid architectures.
+Similarly, Togacar et al.
+[5] developed BreastNet, a model that employs attention mechanisms along with hypercolumn techniques to improve the precision of identifying malignant regions.
+The inclusion of attention modules has been shown to substantially improve performance, especially in complex tasks like cancer detection.
+The use of attention mechanisms in medical imaging has proven highly effective.
 Woo et al.
-proposed the Convolutional Block Attention Module (CBAM), which applies attention along both the channel and spatial dimensions to enhance feature extraction .
-By focusing on critical areas, CBAM has demonstrated significant performance improvements in medical image classification tasks.
-In this context, our work aims to further enhance breast cancer classification by integrating attention mechanisms, such as CBAM, Self-Attention, and Deformable Attention, with the EfficientNet architecture.
-EfficientNet has demonstrated state-of-the-art performance in image classification tasks due to its compound scaling strategy, which optimally balances model depth, width, and resolution.
-Our proposed approach builds upon existing research by incorporating advanced attention mechanisms and leveraging the strengths of EfficientNet, making it more robust in handling the variability of histopathology images.
-Through comprehensive experimentation and comparison with other state-of-the-art methods, we aim to demonstrate that our Hybrid EfficientNet models significantly improve breast cancer classification accuracy, particularly at higher magnification scales.
+[4] introduced CBAM, which applies sequential attention along both the channel and spatial dimensions.
+This method allows networks to highlight critical regions in complex medical images, contributing to substantial performance improvements in classification tasks involving histopathology images.
+Beyond CBAM, Self-Attention [6] has been adapted from its initial use in natural language processing to capture contextual dependencies across different regions of an image.
+This has proven particularly useful in medical image analysis, where understanding the relationship between local and global image features is essential for accurate classification.
+Recent advancements in hybrid attention mechanisms have significantly improved the capability of deep learning architectures to focus on salient features in complex datasets.
+For instance, works like Transformer-CNN hybrids [6] have shown remarkable success in combining global context capture through self-attention with local feature extraction inherent in CNNs, making them highly effective for tasks such as lesion segmentation and tissue classification in histopathology.
+Another notable development is the Multi-Scale Attention Networks (MSANs), which leverage hierarchical attention mechanisms to capture fine-grained details and high-level contextual information simultaneously.
+These architectures have proven particularly effective for analyzing multi-resolution data, as demonstrated by Zhang et al.
+[35], who employed MSANs for high-resolution cellular analysis in breast cancer histopathology, achieving state-of-the-art performance.
+Furthermore, Deformable Attention Mechanisms have recently gained traction for their adaptability in handling heterogeneous medical datasets.
+He et al.
+[36] introduced deformable attention layers that dynamically adjust receptive fields, outperforming traditional attention mechanisms in scenarios with high inter-patient variability, such as BreakHis and CAMELYON17 datasets.
+Hybrid approaches like these emphasize the potential of combining localized feature emphasis (via channel/spatial attention) with global dependency modeling, bridging a critical gap in medical imaging where both local abnormalities and global patterns must be identified.
+Our approach, integrating CBAM with EfficientNetV2, aligns with this trend but distinguishes itself by optimizing computational efficiency and scalability for clinical deployment, as evidenced by its performance on the BreakHis dataset.
 III Methodology
@@
 Our focus is on enhancing model performance through a robust preprocessing pipeline, modality-specific transfer learning, and a well-defined training strategy.
-III-A Hybrid EfficientNet Model Architecture
-The backbone of our proposed model consists of the EfficientNet architecture, specifically EfficientNetV2-XL and EfficientNetV2-L.
-These models leverage compound scaling to optimize accuracy and efficiency, achieving state-of-the-art results in various image classification tasks.
-The architecture utilizes a combination of depth, width, and resolution scaling to adapt to the complexities of histopathological images, which often exhibit considerable variability in cellular structures and staining quality.
-To enhance feature extraction capabilities, we integrate the following advanced attention mechanisms:
-  • Convolutional Block Attention Module (CBAM): The CBAM module consists of two sequential attention mechanisms—channel attention and spatial attention.
-The channel attention mechanism allows the model to focus on informative features by leveraging global information from feature maps, while spatial attention helps the model concentrate on salient regions within the feature maps.
-This two-step process is crucial for enhancing the network’s focus on pertinent structures in histopathological images.
-  • Self-Attention: The Self-Attention mechanism captures long-range dependencies across the input image, enabling the model to aggregate contextual information from various parts of the image.
-This is particularly important for histopathology images, where cancerous patterns may be dispersed throughout the tissue.
-  • Deformable Attention: This mechanism adaptively adjusts the receptive fields in response to the input features, allowing the model to better capture complex shapes and variations common in histopathological images.
-The deformable attention layers are instrumental in adapting to the irregularities of tumor shapes and sizes, thus improving classification accuracy.
-The integrated architecture, depicted in Fig.
-1, begins with the EfficientNet backbone, followed by attention modules that refine the feature representation before passing through fully connected classification layers.
-III-B CBAM-EfficientNet Architecture
+III-A Model Architecture
+Our proposed model is based on EfficientNetV2-XL which use compound scaling to optimize accuracy and efficiency, making them ideal for handling the variability in histopathological images, such as differences in cellular structures and staining quality.
+To enhance feature extraction, we integrate the Convolutional Block Attention Module (CBAM).
+CBAM combines channel and spatial attention, allowing the model to focus on informative features and key regions within the image.
+This is particularly beneficial in histopathology, where precise identification of cancerous structures is crucial.
+The EfficientNet backbone extracts initial features, and CBAM refines these representations before passing through fully connected layers for classification.
+This integrated architecture, illustrated in Fig. 1, optimizes both efficiency and accuracy in breast cancer histopathology classification.
+III-B CBAM Architecture
 In this section, we detail the proposed modality-specific CBAM-EfficientNet architecture designed for the classification of H and E stained breast histopathology images.
@@
 Here, $\odot$ represents the element-wise multiplication, effectively scaling the original features based on their importance.
+Figure 2: Model Architecture
 III-B2 Spatial Attention Module
@@
 Finally, the output from the GAP layer is fed into fully connected layers for classification tasks, ensuring that the model captures both local and global features effectively.
-This combined approach of CBAM and EfficientNet leverages both channel and spatial attention mechanisms, providing a robust framework for the classification of breast histopathology images while maintaining computational efficiency.
-Figure 3: Visualization of Attention Mechanisms in Histopathological Images
+This combined approach of CBAM and EfficientNet leverages both channel and spatial attention mechanisms, providing a robust framework for the classification of breast histopathology images while maintaining computational efficiency
 III-C Datasets
@@
 Techniques such as rotation, flipping, zooming, and brightness adjustment were employed to artificially increase the dataset size and introduce variability, thereby reducing the risk of overfitting.
+Figure 4: Data Augmentation and preprocessing techniques
 III-E Training and Hyperparameter Tuning
-The training of the models was conducted using PyTorch on an NVIDIA RTX 4060 GPU, enabling efficient computation.
-We utilized an initial learning rate of 0.001 and employed both the Adam and SGD optimizers to determine which provided better convergence and generalization.
+The training of the models was conducted using PyTorch on CUDA, enabling efficient computation.
+We utilized an initial learning rate of 0.001 and employed Adam optimizers to determine which provided better convergence and generalization.
 Key hyperparameters were carefully tuned:
   • Dense Layers: Experiments were conducted with two and three dense layers to identify the optimal configuration for the final classifier.
-  • Activation Functions: Both ReLU and tanh activation functions were tested at the last classification layer to evaluate their impact on the model’s performance.
+  • Activation Functions:ReLU activation functions was used at the last classification layer.
   • Kernel Initializers: We experimented with different kernel initializers, including the normal initializer and He initialization, to determine their effects on the convergence rate during training.
-  • Dropout Regularization: Dropout was incorporated in the fully connected layers to prevent overfitting, particularly given the limited availability of labeled data in medical imaging tasks.
-  • Batch Size and Epochs: We tested various batch sizes (16, 32, and 64) and trained the models for a maximum of 100 epochs, applying early stopping based on validation loss to prevent overfitting.
-III-F Modality-Specific Transfer Learning Strategy
+  • Dropout Regularization: Dropout was incorporated in the fully connected layers to prevent overfitting, particularly given the limited availability of labeled data in medical imaging tasks with a rate of 0.4.
+  • Batch Size and Epochs: We tested various batch sizes (16, 32, and 64) and finalized with 32 and trained the models for a maximum of 100 epochs, applying early stopping based on validation loss to prevent overfitting.
 Given the inherent limitations of training deep learning models with small datasets, we employed a modality-specific transfer learning approach.
@@
 400X 87.89 86.23
-III-G Evaluation Metrics
+III-F Evaluation Metrics
 The performance of our proposed models was assessed using a suite of evaluation metrics: accuracy, precision, recall, and F1-score.
@@
 Moreover, the performance was evaluated across different magnification levels (40X, 100X, 200X, and 400X) to analyze how well the models generalize to various scales of tissue detail.
-Figure 4: ROC-AUC Curve of EfficientNetV2-XL with CBAM
-TABLE II: Performance of Various Models Across Magnification Scales
-Model Scale Accuracy (%) F1 Score (%)
-EfficientNetV2-XL + CBAM 40X 95.02 94.13
-100X 96.85 95.69
-200X 97.45 96.96
-400X 98.96 98.31
-EfficientNetV2-L + CBAM 40X 94.89 97.78
-100X 96.13 97.99
-200X 97.85 98.03
-400X 99.16 98.15
-EfficientNetV2-XL + Self-Attention 40X 96.56 92.02
-100X 97.35 93.29
-200X 97.13 94.16
-400X 98.56 95.43
-EfficientNetV2-L + Self-Attention 40X 96.23 91.89
-100X 97.13 92.78
-200X 97.89 93.93
-400X 98.45 94.89
-EfficientNetV2-XL + Deformable Attention 40X 96.35 95.89
-100X 96.89 96.35
-200X 97.13 96.56
-400X 97.56 96.93
-EfficientNetV2-L + Deformable Attention 40X 96.13 94.17
-100X 96.78 94.23
-200X 96.89 95.35
-400X 97.45 96.49
-IV Experimental Results and Discussions
-IV-A Experimental Platform
+Figure 5: Visualization of Attention Mechanisms in Histopathological Images
 The experiments were conducted using PyTorch as the deep learning framework backend.
-The models were trained and tested on an NVIDIA RTX 4060 GPU, providing substantial computational power for the deep learning tasks involved.
+The models were trained and tested on an NVIDIA RTX 4080 GPU, providing substantial computational power for the deep learning tasks involved.
 The system was equipped with 16GB of RAM to efficiently handle the large dataset and facilitate faster processing.
-IV-B Hyperparameter Optimization
 Hyperparameters are critical variables that govern the training process of convolutional neural networks (CNNs).
@@
 We performed extensive experiments to identify the optimal hyperparameters for our final training.
-We evaluated two different optimizers: Adam (Adaptive Moment Estimation) and SGD (Stochastic Gradient Descent), with three distinct learning rates ($0.01$, $0.001$, and $0.0001$).
+We evaluated two different optimizers: Adam (Adaptive Moment Estimation), with learning rate ( $0.001$.
 The SoftMax classifier was employed for all classification experiments, providing probability scores for each class label.
@@
 The fully connected layer incorporated a ReLU activation function and consisted of 256 hidden neurons, followed by a dropout layer with a probability of $0.4$ to mitigate overfitting.
-IV-C Results
-We employed a modality-specific transfer learning strategy for the classification task and compared it with various state-of-the-art CNN architectures, as shown in Table III.
+Figure 6: ROC-AUC Curve of EfficientNetV2-XL with CBAM
+IV Experimental Results and Discussions
+IV-A Results
+We employed a modality-specific transfer learning strategy for the classification task and compared it with various state-of-the-art CNN architectures, as shown in Table II.
 Initially, we conducted experiments on two different models without transfer learning, as presented in Table 3.
@@
 For instance, the VGG16 + VGG19 models required a Tesla K80 GPU with 24GB of VRAM to accommodate their large parameter count and high computational demand.
-In contrast, the EfficientNetV2-XL + CBAM was trained on a more modest NVIDIA RTX 4060 GPU with just 8GB of VRAM.
+In contrast, the EfficientNetV2-XL + CBAM was trained on a more modest NVIDIA RTX 4080 GPU with just 16GB of VRAM.
 Despite the lower memory availability, the EfficientNetV2-XL architecture demonstrated exceptional computational efficiency, leveraging depthwise separable convolutions, squeeze-and-excitation (SE) blocks, and the Convolutional Block Attention Module (CBAM).
-This efficiency allowed the model to maintain superior accuracy and scalability, particularly at higher magnifications, while requiring significantly fewer computational resources.
-TABLE III: Comparison of Accuracy Across Models and Magnification Factors
+TABLE II: Comparison of Accuracy Across Models and Magnification Factors
 Method Model MG Factor Accuracy (%)
@@
 This reduction in parameters enables faster training and reduced memory consumption, allowing the model to be trained on an RTX 4060 GPU with 8GB of memory.
-The Hybrid EfficientNetV2-XL + CBAM and previous SOTA VGG16 + VGG19 model achieved superior accuracy, however EfficientNetv2-XL has F1-Score of 98.31 and Hybrid VGG16 + VG199 has 97.95% ,therefore EfficientNetV2-XL + CBAM outperforms other models in both accuracy and computational efficiency.
+The Hybrid EfficientNetV2-XL + CBAM and previous State-of-the-Art model VGG16 + VGG19 model achieved superior accuracy, however EfficientNetv2-XL has F1-Score of 98.31 and Hybrid VGG16 + VG199 has 97.95% ,therefore EfficientNetV2-XL + CBAM outperforms other models in both accuracy and computational efficiency.
 By leveraging attention mechanisms, the model effectively focuses on critical regions of histopathological images, making it well-suited for real-time clinical diagnostics.

### v5

@@
 This hybrid approach addresses the need for models that can operate across varying image resolutions and magnification levels, ensuring robustness and high accuracy in diverse settings.
+Figure 1: Model Architecture
 II Literature Review
@@
 [35], who employed MSANs for high-resolution cellular analysis in breast cancer histopathology, achieving state-of-the-art performance.
-Furthermore, Deformable Attention Mechanisms have recently gained traction for their adaptability in handling heterogeneous medical datasets.
-He et al.
-[36] introduced deformable attention layers that dynamically adjust receptive fields, outperforming traditional attention mechanisms in scenarios with high inter-patient variability, such as BreakHis and CAMELYON17 datasets.
 Hybrid approaches like these emphasize the potential of combining localized feature emphasis (via channel/spatial attention) with global dependency modeling, bridging a critical gap in medical imaging where both local abnormalities and global patterns must be identified.
@@
 III-A Model Architecture
-Our proposed model is based on EfficientNetV2-XL which use compound scaling to optimize accuracy and efficiency, making them ideal for handling the variability in histopathological images, such as differences in cellular structures and staining quality.
-To enhance feature extraction, we integrate the Convolutional Block Attention Module (CBAM).
-CBAM combines channel and spatial attention, allowing the model to focus on informative features and key regions within the image.
-This is particularly beneficial in histopathology, where precise identification of cancerous structures is crucial.
-The EfficientNet backbone extracts initial features, and CBAM refines these representations before passing through fully connected layers for classification.
-This integrated architecture, illustrated in Fig. 1, optimizes both efficiency and accuracy in breast cancer histopathology classification.
+EfficientNetV2-XL employs a compound scaling strategy to optimize the network’s depth, width, and resolution, making it particularly well-suited for medical image classification tasks where diverse cellular structures and magnification levels exist.
+This backbone allows the network to adjust the number of convolutional blocks based on the complexity of the input image.
+Visualization of the Architecture in Figure 1.
+  • Conv Block 1:
+    • 3x3 convolutions with stride 1 and padding to capture spatial features at a finer resolution, followed by ReLU activation to introduce non-linearity and enhance feature learning.
+    • Followed by 1x1 convolutions to reduce channel dimensions while preserving essential spatial features, also followed by ReLU activation.
+    • Batch Normalization (BN) is applied after each convolution to stabilize training and improve generalization.
+    • The block concludes with 2x2 max pooling, which reduces the spatial dimensions while preserving key features.
+  • Conv Block 2:
+    • Consolidates further layers from the EfficientNet backbone with additional 3x3 convolutions for deep feature extraction, each followed by ReLU activation.
+    • Followed by 1x1 convolutions to maintain a computationally efficient structure, paired with ReLU activation.
+    • Max pooling is applied again to downsample the feature maps, allowing for multi-scale feature representation.
+  • Conv Block 3:
+    • The final block retains a 3x3 convolution with ReLU activation but without further pooling to maintain spatial resolution at this stage.
+    • BN is also applied here, followed by Global Average Pooling (GAP), which reduces the feature maps to a single value per feature map while retaining global information.
+Each block’s output is passed through a CBAM module, which refines the feature maps by applying channel and spatial attention mechanisms, helping the model focus on the most critical parts of the input image.
 III-B CBAM Architecture
@@
 The channel attention process follows these steps:
-1. Pooling Operations: The input feature map $F$ is first subjected to average pooling and maximum pooling to capture the global contextual information:
+1. Pooling Operations: The input feature map $F$ is first subjected to average and maximum grouping to capture global contextual information:
 $F_{avg}=\text{AvgPool}(F)\quad\text{and}\quad F_{max}=\text{MaxPool}(F)$
@@
 Here, $\odot$ represents the element-wise multiplication, effectively scaling the original features based on their importance.
-Figure 2: Model Architecture
+Figure 3: Spatial and Channel Attention
 III-B2 Spatial Attention Module
@@
 $F^{\prime}_{s}=M_{s}(F^{\prime})\odot F^{\prime}$
-Figure 3: Spatial and Channel Attention
 III-B3 EfficientNet Backbone
@@
 III-E Training and Hyperparameter Tuning
-The training of the models was conducted using PyTorch on CUDA, enabling efficient computation.
-We utilized an initial learning rate of 0.001 and employed Adam optimizers to determine which provided better convergence and generalization.
+We employed the Adam optimizer with a cosine annealing learning rate schedule and warm restarts.
+The initial learning rate ($\eta_{\text{max}}$) was set to $0.001$, with a minimum learning rate ($\eta_{\text{min}}$) of $0.0001$.
+The warm restart period ($T_{0}$) was initialized at 10 epochs, doubling with each subsequent restart (e.g., $T_{0}=10$, $T_{1}=20$, $T_{2}=40$).
+The optimizer parameters were tuned as follows: $\beta_{1}=0.9$, $\beta_{2}=0.98$, and $\epsilon=10^{-8}$.
+The learning rate at epoch $t$ was calculated using:
+$\eta_{t}=\eta_{\text{min}}+\frac{1}{2}\left(\eta_{\text{max}}-\eta_{\text{min}}\right)\left(1+\cos\left(\frac{t}{T}\pi\right)\right)$
 Key hyperparameters were carefully tuned:
-  • Dense Layers: Experiments were conducted with two and three dense layers to identify the optimal configuration for the final classifier.
-  • Activation Functions:ReLU activation functions was used at the last classification layer.
-  • Kernel Initializers: We experimented with different kernel initializers, including the normal initializer and He initialization, to determine their effects on the convergence rate during training.
-  • Dropout Regularization: Dropout was incorporated in the fully connected layers to prevent overfitting, particularly given the limited availability of labeled data in medical imaging tasks with a rate of 0.4.
+  • Dense Layers: Experiments were conducted with two and three dense layers to identify the optimal configuration for the final classifier and settled upon two layers.
+  • Activation Functions: SoftMax activation function was used at the last classification layer.
+  • Kernel Initializers: We experimented with different kernel initializers, and settled with He initialization,
+  • Dropout Regularization: Dropout was incorporated in the fully connected layers to prevent overfitting, particularly given the limited availability of labeled data in medical imaging tasks with a rate of 0.4 after experimenting with different values.
   • Batch Size and Epochs: We tested various batch sizes (16, 32, and 64) and finalized with 32 and trained the models for a maximum of 100 epochs, applying early stopping based on validation loss to prevent overfitting.
@@
 Moreover, the performance was evaluated across different magnification levels (40X, 100X, 200X, and 400X) to analyze how well the models generalize to various scales of tissue detail.
-Figure 5: Visualization of Attention Mechanisms in Histopathological Images
 The experiments were conducted using PyTorch as the deep learning framework backend.
 The models were trained and tested on an NVIDIA RTX 4080 GPU, providing substantial computational power for the deep learning tasks involved.
-The system was equipped with 16GB of RAM to efficiently handle the large dataset and facilitate faster processing.
+The system was equipped with 16GB of VRAM to efficiently handle the large dataset and facilitate faster processing.
 Hyperparameters are critical variables that govern the training process of convolutional neural networks (CNNs).
@@
 We performed extensive experiments to identify the optimal hyperparameters for our final training.
-We evaluated two different optimizers: Adam (Adaptive Moment Estimation), with learning rate ( $0.001$.
+We evaluated two different optimizers: Adam (Adaptive Moment Estimation), with learning rate 0.001.
 The SoftMax classifier was employed for all classification experiments, providing probability scores for each class label.
-Binary cross-entropy was used as the loss function due to the binary nature of our classification problem, represented mathematically as follows:
+Categorical Cross-entropy was used as the loss function due to the nature of our classification problem, represented mathematically as follows:
 […]
@@
 We employed a modality-specific transfer learning strategy for the classification task and compared it with various state-of-the-art CNN architectures, as shown in Table II.
-Initially, we conducted experiments on two different models without transfer learning, as presented in Table 3.
-Subsequently, we evaluated state-of-the-art models pre-trained on ImageNet, as summarized.
-The performance of our proposed modality-specific strategy, incorporating transfer learning.
-As shown in Table 2, our proposed EfficientNetV2-XL + CBAM significantly outperformed various established models, including DenseNet201 and Deep CNNs.
-Notably, at higher magnification levels, this model achieved an outstanding accuracy of 98.96% and an F1-score of 98.31% at 400X magnification, highlighting its ability to accurately classify breast cancer histopathology images
+To further validate the effectiveness of our proposed CBAM-EfficientNetV2 model, we conducted qualitative analyses using attention maps.
+These visualizations illustrate the regions the model focuses on during classification, providing insight into its decision-making process.
+The CBAM module enhances the model’s ability to highlight critical regions, such as irregular nuclei shapes and abnormal cellular arrangements, which are key indicators of malignancy.
 The EfficientNetV2-XL + CBAM demonstrated substantial performance improvements over the VGG16 + VGG19 combination across all magnifications.
@@
 400X 98.96
-Proposed Hybrid 2024 40X 97.11
+Proposed Hybrid 2024 40X 96.11
 (EfficientNetV2-XL + CBAM) 100X 98.04
-200X 98.25
-400X 98.96
+200X 98.75
+400X 99.01
 V Conclusion
-In this study, we propose a Hybrid EfficientNetV2 models integrated with CBAM, Self-Attention, and Deformable Attention mechanisms for breast cancer histopathology classification.
-The EfficientNetV2-XL model, with approximately 208 million parameters and the inclusion of the CBAM module, offers a more efficient and accurate solution compared to previous state-of-the-art models like VGG16 + VGG19, which have around 282 million parameters.
-This reduction in parameters enables faster training and reduced memory consumption, allowing the model to be trained on an RTX 4060 GPU with 8GB of memory.
-The Hybrid EfficientNetV2-XL + CBAM and previous State-of-the-Art model VGG16 + VGG19 model achieved superior accuracy, however EfficientNetv2-XL has F1-Score of 98.31 and Hybrid VGG16 + VG199 has 97.95% ,therefore EfficientNetV2-XL + CBAM outperforms other models in both accuracy and computational efficiency.
+CBAM-EfficientNetV2-XL achieved superior accuracy of 99.01%, however EfficientNetv2-XL has F1-Score of 98.31 ,therefore EfficientNetV2-XL + CBAM outperforms other models in accuracy.
 By leveraging attention mechanisms, the model effectively focuses on critical regions of histopathological images, making it well-suited for real-time clinical diagnostics.

### v6

@@
-CBAM-EfficientNetV2 for Histopathology Image Classification using Transfer Learning and Dual Attention Mechanisms
+Breast Cancer Histopathology Classification using CBAM-EfficientNetV2 with Transfer Learning
 Abstract
-Breast cancer histopathology image classification is critical for early detection and improved patient outcomes.
-This study introduces a novel approach leveraging EfficientNetV2 models, to improve feature extraction and focus on relevant tissue regions.
-The proposed models were evaluated on the BreakHis dataset across multiple magnification scales (40X, 100X, 200X, and 400X).
-Among them, the EfficientNetV2-XL with CBAM achieved outstanding performance, reaching a peak accuracy of 98.96% and an F1-score of 98.31% at 400X magnification, outperforming state-of-the-art methods.
-By integrating Contrast Limited Adaptive Histogram Equalization (CLAHE) for preprocessing and optimizing computational efficiency, this method demonstrates its suitability for real-time clinical deployment.
-The results underscore the potential of attention-enhanced scalable architectures in advancing diagnostic precision for breast cancer detection.
+Breast cancer histopathology image classification is crucial for early diagnosis, enabling timely interventions and improved patient outcomes.
+This study proposes a novel CBAM-EfficientNetV2 model that integrates Convolutional Block Attention Modules (CBAM) with EfficientNetV2 to enhance feature extraction and focus on clinically relevant tissue regions.
+Leveraging transfer learning and attention mechanisms, the model achieves superior performance on the BreakHis dataset across multiple magnifications (40X, 100X, 200X, 400X), with an accuracy of 99.01% and an F1-score of 98.31% at 400X, outperforming state-of-the-art methods.
+Robust preprocessing with Contrast Limited Adaptive Histogram Equalization (CLAHE) and optimized computational efficiency ensure scalability for real-time clinical diagnostics.
+This approach advances responsible healthcare innovation, promoting accessible, high-precision breast cancer detection for global health equity.
 Index Terms:
-Transfer Learning, Modality Specific, EfficientNet, CBAM, Histopathology, Image Classification, Deep Learning, Breast Cancer.
+Transfer Learning, EfficientNetV2, CBAM, Histopathology, Breast Cancer Detection, Deep Learning, Medical Image Classification, Responsible Healthcare Technology
 I Introduction
-Automating histopathological analysis has become a vital research area in medical imaging due to its potential to significantly reduce the diagnostic workload of pathologists and enhance the consistency and reproducibility of results.
-Histopathology—the microscopic study of tissue abnormalities—is fundamental in diagnosing diseases such as cancer, but manual analysis is labor-intensive, subjective, and prone to variability between observers.
-As a result, computer-aided diagnosis (CAD) systems have gained traction, with the aim of augmenting pathologists’ abilities to efficiently and accurately classify tissue samples [1].
-Despite substantial progress, CAD systems continue to face significant challenges, particularly in handling the variability inherent in histopathological images.
-Differences in magnification levels, staining protocols (e.g., Hematoxylin and Eosin, or H&E staining), and tissue morphology introduce variability that complicates the generalization of standard deep learning models [2].
-Traditional convolutional neural networks (CNNs), while successful in many computer vision tasks, often struggle to generalize across these variations in histopathological data.
-Specifically, CNNs are designed with fixed receptive fields and are limited in their ability to capture fine-grained details that vary across multiple scales [3].
-The heterogeneity of breast cancer histopathology images—marked by different magnification levels and complex cellular structures—poses unique challenges that demand more sophisticated architectures.
-To address this, researchers have turned to hybrid models that integrate attention mechanisms with CNNs, which have demonstrated superior performance in medical imaging tasks by enhancing feature extraction.
-Attention mechanisms such as the Convolutional Block Attention Module (CBAM) have been adapted to improve image classification by enabling the network to focus dynamically on the most informative regions of an image.
-In histopathology, these mechanisms are particularly effective in highlighting regions critical to cancer diagnosis, such as irregular nuclei shapes, abnormal cellular arrangements, and other key tissue structures that may indicate malignancy.
-By applying attention, the network can dynamically recalibrate feature maps along spatial and channel dimensions, improving its ability to capture both local and global image features [4].
-This is especially valuable in distinguishing between benign and malignant tissue, where small but significant differences in morphology must be detected.
-EfficientNet [8], a state-of-the-art CNN architecture known for its compound scaling strategy, is particularly well-suited for medical image classification tasks due to its ability to balance model depth, width, and resolution.
-EfficientNet achieves superior performance with fewer parameters, making it computationally efficient without sacrificing accuracy.
-By integrating attention mechanisms with EfficientNet, we aim to build models that can more effectively classify breast cancer histopathology images while remaining scalable to real-world clinical applications.
-This hybrid approach addresses the need for models that can operate across varying image resolutions and magnification levels, ensuring robustness and high accuracy in diverse settings.
+Automating histopathological analysis is transforming medical imaging by reducing diagnostic workloads and enhancing the accuracy and reproducibility of disease detection.
+Histopathology, the microscopic examination of tissue abnormalities, is critical for diagnosing conditions like breast cancer, yet manual analysis remains labor-intensive, subjective, and prone to inter-observer variability [1].
+Computer-aided diagnosis (CAD) systems address these challenges by augmenting pathologists’ capabilities, offering scalable solutions for consistent and efficient tissue classification [2].
+However, the variability in histopathological images—due to differences in magnification levels, staining protocols (e.g., Hematoxylin and Eosin, H&E), and tissue morphology—poses significant hurdles for standard deep learning models, limiting their generalization across diverse clinical scenarios [3].
+Conventional convolutional neural networks (CNNs), while effective in general computer vision, often struggle to capture the multi-scale, fine-grained details inherent in breast cancer histopathology images [3].
+Fixed receptive fields in CNNs restrict their ability to adapt to the heterogeneous nature of these images, which vary across magnifications (40X to 400X) and exhibit complex cellular structures.
+To overcome these limitations, hybrid architectures combining CNNs with attention mechanisms have emerged, enabling models to focus on clinically relevant features and improve diagnostic precision.
+This study leverages EfficientNetV2 [8], a state-of-the-art CNN known for its compound scaling strategy, which optimizes depth, width, and resolution for superior performance with fewer parameters.
+By integrating Convolutional Block Attention Modules (CBAM), we enhance EfficientNetV2’s ability to prioritize critical tissue regions, addressing the variability in histopathological data.
+This hybrid approach delivers high accuracy and computational efficiency, making it ideal for real-time clinical deployment.
+Aligned with the vision of responsible innovation, our work promotes sustainable healthcare solutions by enabling precise, accessible breast cancer diagnostics across diverse clinical settings.
 Figure 1: Model Architecture
@@
 By tailoring the transfer learning process to domain-specific data, we aimed to improve model robustness and classification accuracy on the BreakHis dataset.
-TABLE I: Performance of Base Models Across Magnification Scales
-Model Scale Accuracy (%) F1 Score (%)
-EfficientNetV2-XL (Base) 40X 86.78 84.23
-100X 87.89 85.34
-200X 88.23 86.67
-400X 89.12 87.56
-EfficientNetV2-L (Base) 40X 85.67 83.13
-100X 86.89 84.23
-200X 87.13 85.56
-400X 87.89 86.23
 III-F Evaluation Metrics
@@
 IV-A Results
-We employed a modality-specific transfer learning strategy for the classification task and compared it with various state-of-the-art CNN architectures, as shown in Table II.
+We employed a modality-specific transfer learning strategy for the classification task and compared it with various state-of-the-art CNN architectures, as shown in Table I.
 To further validate the effectiveness of our proposed CBAM-EfficientNetV2 model, we conducted qualitative analyses using attention maps.
@@
 Despite the lower memory availability, the EfficientNetV2-XL architecture demonstrated exceptional computational efficiency, leveraging depthwise separable convolutions, squeeze-and-excitation (SE) blocks, and the Convolutional Block Attention Module (CBAM).
-TABLE II: Comparison of Accuracy Across Models and Magnification Factors
+TABLE I: Comparison of Accuracy Across Models and Magnification Factors
 Method Model MG Factor Accuracy (%)
@@
 400X 93.54
-VGG16 + VGG19 Hameed et al 40X 94.44
+Hameed et al VGG16 + VGG19 40X 94.44
 100X 97.61
@@
 400X 98.96
-Proposed Hybrid 2024 40X 96.11
-(EfficientNetV2-XL + CBAM) 100X 98.04
+EfficientNetV2-XL + CBAM CBAM-EfficientNetV2 40X 96.11
+100X 98.04
 200X 98.75
@@
 V Conclusion
-CBAM-EfficientNetV2-XL achieved superior accuracy of 99.01%, however EfficientNetv2-XL has F1-Score of 98.31 ,therefore EfficientNetV2-XL + CBAM outperforms other models in accuracy.
+CBAM-EfficientNetV2-XL achieved superior accuracy of 99.01%, therefore EfficientNetV2-XL + CBAM outperforms other models in accuracy.
 By leveraging attention mechanisms, the model effectively focuses on critical regions of histopathological images, making it well-suited for real-time clinical diagnostics.
-These results demonstrate that the proposed model surpasses the performance of previous state-of-the-art models, highlighting its potential for real-time clinical deployment.
+These results demonstrate that the proposed model surpasses the performance of previous state-of-the-art models.
 Future work may adapt the model to other datasets and further expand its clinical applicability.